# M2SNet: Multi-scale in Multi-scale Subtraction Network for Medical Image Segmentation

### v2

@@ 2.1 Medical Image Segmentation Network
According to the characteristics of different organs or lesions, we classify existing medical image segmentation methods into two types: medical-general and medical-specific one.
Medicine-general Methods. With the U-Net UNet achieving stable performance in the medical image segmentation field, the U-shape structure with encoder-decoder has become the basic segmentation baseline. U-Net++ UNet++ integrates both the long connection and short connection, which can reduce the semantic gap between the feature maps of the encoder and decoder sub-networks. For attention U-Net Attention_UNet , an attention gate is embedded in each transition layer between the encoder and decoder block, which can automatically learn to focus on target structures of varying shapes and sizes. More recently, Transformer-based approaches Segmenter ; SegFormer ; SwinUNet have gained prominence by exploiting self-attention to capture long-range dependencies beyond the limited receptive fields of CNNs. SegFormer SegFormer and SwinUNet SwinUNet combine hierarchical feature extraction with attention-driven decoding, while hybrid designs such as UTNet UTNet and TransUNet TransUNet_MIA embed Transformer modules into U-shaped frameworks to infuse global context into local representations. Despite their improved expressiveness, these models often suffer from quadratic complexity and limited local inductive bias, which can hinder scalability and fine-grained predictions. In response, state-space–based architectures have emerged as an efficient alternative. Methods built on Mamba Mamba , such as Sigma Sigma , SegMamba Segmamba and U-Mamba U-mamba , replace attention with structured recurrence, achieving linear time complexity and reduced memory footprint. By combining the inherent structural bias of state-space models with CNNs’ local detail preservation, hybrid CNN–Mamba designs offer a promising route to scalable, high-resolution, and real-time medical image segmentation.
Medicine-specific Methods. In the polyp segmentation task, SFA SFA and PraNet PraNet , focus on recovering the sharp boundary between a polyp and its surrounding mucosa. The former proposes a selective feature aggregation structure and a boundary-sensitive loss function under a shared encoder and two mutually constrained decoders. The latter utilizes a reverse attention module to establish the relationship between the region and boundary cues. UM-Net UM-Net introduces color transfer to reduce color–polyp dependency and incorporating uncertainty estimation with variance correction to enhance the reliability of segmentation results. In addition, Ji et al. PNSNet utilize spatio-temporal information to build the video polyp segmentation model. In the lung infection task, Paluru et al. Anam-Net propose an anamorphic depth embedding-based lightweight CNN to segment anomalies in chest CT images. Inf-Net Inf-Net builds the implicit reverse attention and explicit edge attention to model the boundaries. BCS-Net BCS-Net has three progressive boundary context-semantic reconstruction blocks, which can help the decoder to capture the piecemeal region for lung infection. In the breast segmentation task, Byra et al. SKUNet develop a selective kernel via an attention mechanism to adjust the receptive fields of the U-Net, which can further improve the segmentation accuracy of breast tumors. Chen et al. NU-net propose a nested U-net to achieve robust representation of breast tumors by exploiting different depths and sharing weights.
@@ 2.2 Multi-scale Feature Extraction
Scale cues play an important role in capturing contextual information of objects. Inspired by the scale-space theory that has been widely validated as an effective and theoretically sound framework, more and more multi-scale methods are proposed. Compared with single-scale features, multi-scale features are beneficial to address naturally occurring scale variations. This characteristic can help the medical segmentation models perceive lesions with different scales. According to the form, current multi-scale based methods can be roughly divided into two categories, namely, the inter-layer multi-scale structure and the intra-layer multi-scale structure. Inter-layer multi-scale structure is based on features with different scales extracted by the feature encoder and progressively aggregates them in decoder, such as the U-shape UNet ; UNet++ ; FPN ; ACSNet ; PraNet ; U2Net ; DSS ; MINet architecture. Among them, dense skip connections are widely used in decoder because of the advantages in gradient backpropagation and information aggregation. U-Net++ UNet++ undergoes a dense convolution block and re-designs skip pathways to transform the connectivity of the encoder and decoder sub-networks. ICUnet++ ICUnet++ replaces the convolutional layer of U-Net++ with inception structure and add attention gate module before each skip connection to filter interference information. DSSNet DSSNet and CPFP CPFP_RGBDSOD use a fluid pyramid integration strategy to make better use of multi-scale cross-modal/level features. MINet MINet proposes the self-interaction and aggregate interaction strategy to avoid the interference in feature fusion caused by large resolution differences. Intra-layer multi-scale structure usually equips the multi-scale pluggable modules, such as ASPP ASPP , DenseASPP DenseASPP , FoldASPP GateNetv2 , and PAFEM DANet_RGBDSOD to construct the parallel multi-branch convolution layers with different dilated rates to obtain a rich combination of receptive fields. In this work, we apply our subtraction unit to both inter-layer and intra-layer multi-scale feature fusion, which can fully show the flexibility and effectiveness of the subtraction unit, and its general gains will be more convincing. It should be emphasized that we naturally aggregate multi-scale subtractive features in the decoder without focusing on skip connection designs.
@@
2.3 Loss Method
@@ 4.4 Comparisons with Medicine-general and Medicine-specific Methods
Table 8: Quantitative comparison of unified models on three different medical lesion segmentation tasks. The best results are shown in red.
@@
UniverSeg UniverSeg ResNet-101 ResNet 0.673 0.368 0.775 0.600 0.553 0.261
@@
4.5 Comparisons with Unified and Generalist Methods
4.6 Ablation Study
@@ 4.6.1 Effectiveness of the subtraction unit, inter-layer multi-scale subtraction aggregation
4.6.2 Effectiveness of loss function
In Tab. 10, we thoroughly verify the effectiveness of the loss function used in M2SNet. The gap between M2SNet and M2SNet (w/o $\mathcal{L}_{f}$ ) demonstrates the general effectiveness of LossNet for different medical segmentation tasks. At the same time, without the auxiliary of LossNet, the performance of M2SNet itself is good enough to surpass most of the methods in Tab. 1 - Tab. 4. The gap between M2SNet and M2SNet (w/o $\mathcal{L}_{IoU}^{w}$ + $\mathcal{L}_{BCE}^{w}$) shows the necessity of the weighted IoU and BCE loss, which provide a basic foreground region guidance for LossNet to focus on supervising multi-level lesion regions without distracting in the background area.
@@ 4.6.3 Effectiveness of the intra-layer multi-scale subtraction design
Compared to the previous MSNet, the M2SNet replace all the original single-scale subtraction unit with the stronger multi-scale subtraction unit. As shown in Tab. 11, more scales of feature fusion can help improve the overall performance and multi-scale filters of size $[1,3,5]$ and $[1,3,5,7]$ have close performance. Next, we replace the fixed full one weights with Gaussian weights. It can be seen that Gaussian weights significantly degrade the performance of the model in all tasks. We think that the subtraction unit should try to maintain the characteristic distribution of the input features itself, but the Gaussian weights rigidly changes the spatial distribution of the original features, causing an extra burden for the subsequent decoder. Therefore, we choose the multi-scale convolution filters with fixed full one weights of size $[1,3,5]$ as the final setting. To further show the advantages of our intra-layer multi-scale design, we apply other popular multi-scale modules (i.e., ASPP ASPP and DenseASPP DenseASPP ) to the subtraction unit and these architectures are shown in Fig. 7. In Tab. 12, we thoroughly compare both the efficiency and accuracy of these three structures. It can be seen that “M2SNet (Ours)” has a significant performance gain in terms of fourteen metrics on four challenges datasets under different tasks. However, the other two models not only increase the computational burden by more than 50%, but also produce negative gains in multiple datasets. Therefore, the proposed intra-layer multi-scale subtraction design can be taken as a new baseline for future research in subtraction family.
@@
5 Discussion

### v3

@@ 2.1 Medical Image Segmentation Network
Existing medical image segmentation methods can be broadly categorized into medical-general and medical-specific approaches.
Medicine-general Methods. Since the success of U-Net UNet , encoder–decoder architectures with skip connections have become the standard paradigm. Variants such as U-Net++ UNet++ and Attention U-Net Attention_UNet improve feature aggregation and target localization through redesigned skip pathways or attention mechanisms. More recently, Transformer-based models SegFormer ; SwinUNet ; UTNet ; TransUNet_MIA introduce self-attention to capture long-range dependencies, enhancing global context modeling. However, their quadratic complexity and relatively weak local inductive bias may limit scalability and fine-grained prediction. To address efficiency issues, state-space–based architectures such as Mamba Mamba and its segmentation variants Sigma ; Segmamba ; U-mamba replace attention with structured recurrence, achieving linear complexity while maintaining long-range modeling. Hybrid CNN–Mamba frameworks further balance global context modeling and local detail preservation.
Medicine-specific Methods. Task-oriented designs tailor architectures to specific organs or lesions. For example, polyp segmentation models SFA ; PraNet ; UM-Net ; PNSNet emphasize boundary refinement and contextual modeling. Lung infection segmentation methods Anam-Net ; Inf-Net ; BCS-Net focus on edge modeling and anomaly localization, while breast tumor segmentation approaches SKUNet ; NU-net enhance receptive field adaptation and representation robustness. Overall, medical-general methods target universal challenges such as multi-scale representation and cross-level feature aggregation, whereas medical-specific models introduce task-dependent modules, including attention mechanisms and boundary enhancement strategies. Nevertheless, most existing approaches rely heavily on addition or concatenation for feature fusion, which may dilute discriminative difference information. In contrast, our multi-scale subtraction module explicitly models feature discrepancies, enabling more targeted and efficient representation learning.
@@ 2.2 Multi-scale Feature Extraction
Multi-scale modeling is essential for handling objects with large scale variations in medical images. Existing strategies can be grouped into inter-layer and intra-layer designs. Inter-layer multi-scale structures aggregate features from different encoder stages within U-shaped frameworks UNet ; UNet++ ; FPN ; PraNet ; U2Net . These architectures progressively fuse hierarchical features to combine semantic and spatial information. Intra-layer multi-scale structures employ parallel convolutional branches with different receptive fields, such as ASPP ASPP , DenseASPP DenseASPP , and related variants, to capture contextual information at multiple scales. Different from prior work that focuses primarily on feature enrichment via aggregation, we incorporate subtraction operations into both inter-layer and intra-layer multi-scale fusion. By explicitly extracting difference information rather than merely accumulating features, our design enhances discriminative representation while maintaining structural simplicity.
@@
2.3 Segmentation Loss Functions and Perceptual Supervision
@@ 4.4 Comparisons with Medicine-general and Medicine-specific Methods
Table 8: Quantitative comparison of medicine-specific methods on four different medical lesion segmentation tasks in terms of ASD and HD95.
ETIS CVC-300-TV Lung Infection BUSI
Methods ASD($m$) $\downarrow$ HD95($m$) $\downarrow$ ASD($m$) $\downarrow$ HD95($m$) $\downarrow$ ASD($m$) $\downarrow$ HD95($m$) $\downarrow$ ASD($m$) $\downarrow$ HD95($m$) $\downarrow$
PraNet 1.154 2.230 - - - - - -
PNS-Net - - 1.726 3.074 - - - -
Inf-Net - - - - 4.589 9.845 - -
NU-Net - - - - - - 3.576 8.745
MSNet 0.602 1.544 0.904 1.996 2.322 8.074 1.967 7.024
M2SNet 0.584 1.280 0.847 1.874 2.021 7.412 1.819 6.451
Table 9: Quantitative comparison of unified, generalist and Mamba-based models on three different medical lesion segmentation tasks. The best results are shown in red.
@@
SegMamba Segmamba Customized Design 0.680 0.802 0.827 0.848 0.815 0.848
H-vmunet H-vmunet Customized Design 0.684 0.809 0.818 0.840 0.819 0.851
UniverSeg UniMRSeg ResNet-101 ResNet 0.673 0.368 0.775 0.600 0.553 0.261
@@
4.5 Evaluation on Boundary Accuracy
In addition to overlap-based metrics such as Dice and IoU, distance-based metrics are widely used to evaluate boundary accuracy in medical image segmentation. In particular, the Average Surface Distance (ASD) measures the average symmetric distance between predicted and ground-truth surfaces, while the 95th percentile Hausdorff Distance (HD95) captures the worst-case boundary discrepancy under a robust percentile criterion. Compared with region-overlap metrics, these indicators provide more direct assessment of contour precision and structural consistency. We conduct comparison with medicine-specific methods across four different tasks. As shown in Tab. 8, M2SNet consistently achieves the lowest ASD and HD95 scores across all tasks, indicating more accurate boundary delineation and reduced extreme surface deviations. These improvements further validate the effectiveness of the proposed multi-scale subtraction design in preserving fine-grained structural information. Importantly, the superiority observed under overlap-based metrics remains consistent when evaluated using distance-based criteria, demonstrating the robustness and reliability of the proposed framework.
4.6 Comparisons with Unified, Generalist and Mamba-based Methods
In addition to unified frameworks, we further compare with recent Mamba-based and SAM-based medical segmentation methods. Specifically, SegMamba Segmamba and H-vmunet H-vmunet represent state-space–based architectures tailored for medical imaging, while MaskSAM MaskSAM and MedSAM-2 MedSAM-2 adapt foundation segmentation models to the medical domain. As shown in Tab. 9, although SegMamba and H-vmunet demonstrate competitive performance compared with some unified generalist models, their overall results remain consistently lower than those of M2SNet across all three tasks and evaluation metrics. This indicates that while state-space modeling improves long-range dependency capture and computational efficiency, it does not inherently guarantee superior lesion discrimination capability in cross-modality scenarios. Similarly, MaskSAM and MedSAM-2 benefit from strong pre-training on large-scale data and exhibit relatively stable performance. However, their segmentation accuracy is still inferior to M2SNet, especially in scenarios involving subtle boundary variations and heterogeneous lesion appearances. This suggests that directly adapting foundation segmentation models to medical imaging, even with domain-specific fine-tuning, may not fully address the strong context dependence and fine-grained structural requirements of lesion segmentation. From a modeling perspective, Mamba-based architectures primarily emphasize efficient sequence modeling and long-range dependency propagation. However, medical lesion segmentation often relies more heavily on precise boundary modeling, multi-scale structural aggregation, and strong local inductive bias. Pure state-space modeling lacks the explicit spatial hierarchy and strong locality priors that are naturally embedded in convolutional designs. Although hybrid variants partially alleviate this limitation, their representation learning capacity for fine-grained lesion structures remains constrained. As a result, CNN-based and Transformer-based architectures continue to dominate mainstream medical image segmentation research.
@@ 4.6 Comparisons with Unified, Generalist and Mamba-based Methods
[…]
Table 12: Ablation experiments on different pre-trained networks used as LossNet.
ETIS CVC-300-TV Lung Infection BUSI
[…]
@@ 4.7.1 Effectiveness of the subtraction unit, inter-layer multi-scale subtraction aggregation
4.7.2 Effectiveness of loss function
In Tab. 11, we thoroughly verify the effectiveness of the loss function used in M2SNet. The gap between M2SNet (LossNet) and M2SNet (w/o LossNet) demonstrates the general effectiveness of LossNet for different medical segmentation tasks. At the same time, without the auxiliary of LossNet, the performance of M2SNet itself is good enough to surpass most of the methods in Tab. 1 - Tab. 4. The gap between M2SNet (LossNet) and M2SNet (w/o $\mathcal{L}_{IoU}^{w}$ & $\mathcal{L}_{BCE}^{w}$) shows the necessity of the weighted IoU and BCE loss, which provide a basic foreground region guidance for LossNet to focus on supervising multi-level lesion regions without distracting in the background area.
We further verify the effectiveness of LossNet from the perspective of supervision strategy. Applying losses at different decoder scales, commonly referred to as deep supervision, mainly enforces pixel-level consistency between intermediate predictions and ground-truth masks. Although deep supervision can accelerate convergence and stabilize gradient propagation, it operates directly in the prediction space and typically relies on carefully designed task-specific loss functions. In contrast, LossNet introduces supervision in a representation space rather than directly in the prediction space. By leveraging a VGG-16 network pre-trained on ImageNet, LossNet extracts hierarchical feature maps that encode rich visual cues, including fine-grained textures, structural patterns, and global semantic context. These multi-level perceptual representations provide comprehensive guidance, encouraging the predicted masks to match the structural and semantic characteristics of the ground truth from a higher-level perspective. Moreover, the rich representations extracted by the pre-trained VGG-16 allow LossNet to employ a simple L2 loss for supervision, reducing dependence on manually designed segmentation-specific losses. To further examine the relationship between LossNet and deep supervision, we conduct additional ablation experiments in Tab. 11. The results show that LossNet complements deep supervision and consistently improves segmentation performance, whereas applying a simple L2 loss directly at decoder scales does not yield comparable gains. This validates the effectiveness of the proposed feature-level perceptual supervision.
4.7.3 Impact of Different Pre-trained Backbones in LossNet
We compare different pre-trained networks as the LossNet, as shown in Tab. 12. The results show that both ResNet-50 ResNet and Swin-T Swin perform worse than VGG-16 VGG when used as perceptual supervision. In fact, in perceptual loss based supervision, a relatively stable community consensus has been established, where most prior works Ploss1 ; Ploss2 ; Esrgan prefer to adopt VGG-16 or VGG-19 as pre-trained feature extractors. The primary reason is that VGG networks provide dense and well-structured hierarchical feature representations, which effectively capture low-level textures, mid-level structural patterns, and high-level semantic information. These multi-level perceptual cues are particularly suitable for segmentation tasks that require accurate modeling of boundary details and global structural consistency. In comparison, ResNet introduces residual connections that enhance semantic abstraction but often lead to more mixed feature representations across layers. This characteristic may reduce the interpretability and stability of feature matching when used for perceptual supervision. Meanwhile, transformer-based architectures emphasize global context modeling through self-attention mechanisms. Although they provide strong semantic representation capability, their token-based feature representation is generally less sensitive to fine-grained spatial textures and boundary details, which are crucial for dense prediction tasks.
@@ 4.7.4 Effectiveness of the intra-layer multi-scale subtraction design
Compared to the previous MSNet, the M2SNet replace all the original single-scale subtraction unit with the stronger multi-scale subtraction unit. As shown in Tab. 13, more scales of feature fusion can help improve the overall performance and multi-scale filters of size $[1,3,5]$ and $[1,3,5,7]$ have close performance. This indicates that incorporating multi-scale difference modeling is beneficial, while excessively large receptive fields do not necessarily provide additional gains. The selection of kernel sizes $1\times 1$, $3\times 3$, and $5\times 5$ is motivated by balancing local detail preservation and moderate receptive field expansion. Specifically, the $1\times 1$ kernel captures channel-wise interactions and preserves fine-grained spatial consistency. The $3\times 3$ kernel models local structural patterns and boundary information, while the $5\times 5$ kernel further enlarges the receptive field to encode mid-range contextual dependencies. These three scales jointly provide complementary local-to-mid-range spatial coverage without introducing excessive computational overhead. We further investigate alternative kernel combinations, including $[1,3,7]$ and $[1,5,9]$, as reported in Tab. 13. The results show that larger-scale combinations do not yield consistent performance improvements compared with $[1,3,5]$, while increasing computational complexity. This suggests that deeper convolutional stacking already provides sufficient receptive field expansion, and overly large kernels may introduce redundant smoothing that weakens boundary modeling. Next, we replace the fixed full one weights with Gaussian weights. It can be seen that Gaussian weights significantly degrade the performance of the model in all tasks. We think that the subtraction unit should try to maintain the characteristic distribution of the input features itself, but the Gaussian weights rigidly changes the spatial distribution of the original features, causing an extra burden for the subsequent decoder. Therefore, we choose the multi-scale convolution filters with fixed full one weights of size $[1,3,5]$ as the final setting. To further show the advantages of our intra-layer multi-scale design, we apply other popular multi-scale modules (i.e., ASPP ASPP and DenseASPP DenseASPP ) to the subtraction unit and these architectures are shown in Fig. 7. In Tab. 14, we thoroughly compare both the efficiency and accuracy of these three structures. It can be seen that “M2SNet (Ours)” has a significant performance gain in terms of fourteen metrics on four challenges datasets under different tasks. However, the other two models not only increase the computational burden by more than 50%, but also produce negative gains in multiple datasets. Therefore, the proposed intra-layer multi-scale subtraction design can be taken as a new baseline for future research in subtraction family.
@@
Table 15: Ablation experiments on different backbone architectures.
ETIS CVC-300-TV Lung Infection BUSI
[…]
4.7.5 Effectiveness of using different backbones
The motivation for using Res2Net-50 in this paper for extensive comparison is as follows: (1) Stronger Multi-scale Feature Representation. Res2Net introduces hierarchical residual-like connections within each residual block by splitting feature channels into multiple groups. This design enables the backbone to capture multi-scale receptive fields inside a single block, allowing the model to simultaneously learn fine-grained local details and long-range contextual information. Such capability is particularly beneficial for medical image segmentation, where target structures often present large scale variations. (2) Better Compatibility with the Proposed Multi-scale Interaction Strategy. Our framework relies on multi-scale in multi-scale subtraction operations to model complementary information across hierarchical feature levels. The enhanced intra-block multi-scale representations provided by Res2Net produce richer scale-aware features, which help the subtraction mechanism better highlight structural differences between feature levels, thereby improving boundary localization and structural consistency. (3) Fair Comparison and Methodological Continuity. The earlier conference version of this work focuses on polyp segmentation, where most state of the art polyp segmentation methods adopt Res2Net as the backbone. To ensure fair comparison with existing approaches and maintain methodological continuity between the conference version and the extended manuscript, we consistently adopt Res2Net-50 in this work.
We further conduct experiments using a stronger CNN-based backbone, ConvNeXt-B ConvNeXt , and a transformer-based backbone, PVTv2-B4 PVTv2 . As shown in Tab. 15, the segmentation performance is further improved with both backbones. These results demonstrate that the proposed M2SNet is compatible with more advanced backbone architectures and exhibits strong generalization capability across different network paradigms.
5 Discussion and Future Work
Subtraction Operation: We provide a theoretical interpretation to explain why subtraction can effectively reduce feature redundancy compared with addition and concatenation. (1) Feature Fusion from Linear Signal Decomposition Perspective. Let $F_{l}$ and $F_{h}$ denote two hierarchical feature representations extracted from adjacent network levels. These features can be decomposed into shared components and complementary components:
$F_{l}=S+C_{l},\quad F_{h}=S+C_{h}$ (7)
where $S$ represents shared semantic information and $C_{l},C_{h}$ denote scale-specific complementary information. Under this formulation, conventional addition-based fusion can be expressed as:
$F_{add}=F_{l}+F_{h}=2S+C_{l}+C_{h}.$ (8)
This formulation preserves and amplifies the shared information $S$, which increases feature redundancy and may cause the fused representation to be dominated by duplicated semantic responses. Similarly, concatenation-based fusion can be written as:
$F_{cat}=[F_{l},F_{h}].$ (9)
This operation retains all feature components, including duplicated shared signals. As a result, the network must rely on additional convolution layers to implicitly suppress redundancy, which increases parameter complexity and computational overhead. In contrast, subtraction-based fusion can be formulated as:
$F_{sub}=F_{h}-F_{l}=C_{h}-C_{l}.$ (10)
This formulation explicitly cancels the shared component $S$, enabling subtraction to naturally suppress redundant information by isolating complementary structural signals between hierarchical features. (2) Correlation Suppression and Orthogonalization Effect. Hierarchical feature maps in encoder–decoder architectures are typically highly correlated due to progressive feature abstraction. From a statistical perspective, addition and concatenation preserve correlated feature components, while subtraction acts as a decorrelation operator. By suppressing correlated shared responses, subtraction encourages the network to learn representations that are closer to complementary or orthogonal feature subspaces, improving representation diversity and discriminability. (3) Frequency Domain Interpretation. Hierarchical features often encode semantic information at different spatial frequency ranges. Lower-level features usually contain high-frequency structural details, while higher-level features capture low-frequency semantic context. Subtraction behaves similarly to a differential operator that enhances structural transitions and high-frequency residual components. This property helps emphasize boundary details and fine structures that are crucial for precise segmentation.
@@ 5 Discussion and Future Work
Limitation: Despite these advantages, several limitations remain. First, the current framework is fully supervised and relies on sufficient annotated training data. In extremely low-data or unseen clinical scenarios, the generalization ability may still be constrained. Second, the multi-scale subtraction unit adopts fixed kernel configurations, which, although computationally efficient and empirically effective, may not fully capture very long-range dependencies compared to dynamic convolution or attention-based mechanisms. Third, while LossNet enhances representation-level supervision, it introduces additional computational overhead during training, which may limit scalability for very large models or high-resolution volumetric inputs.
Future Work: Future research can be explored in several directions. One promising direction is to design adaptive or dynamic multi-scale subtraction mechanisms that can automatically adjust receptive fields according to lesion size and modality characteristics. Another direction is to investigate lightweight perceptual supervision strategies or self-supervised representation alignment methods to reduce training cost while maintaining structural consistency. Furthermore, integrating foundation-model-based pretraining or prompt-driven adaptation may improve generalization under limited annotations. Finally, extending the proposed framework to 3D volumetric segmentation and multi-modal medical imaging scenarios could further enhance its practical applicability in real clinical settings.